# Asking Fast Radio Bursts (FRBs) for More than Reionization History

### v1

@@ 2.2 Structure Function
Figure 1: Image slice from the LC box of the fiducial EoR toy model. Both axes are in comoving units with the axis labelled $x$ being the LoS axis. The bubble radius is $10$ grid units ($\approx 12~{}{\rm Mpc}$). Black and white regions are respectively completely ionized and completely neutral. Reionization progresses from right to left in this box.
We aim to outline how FRBs can be used as a probe for the characteristic size of the ionization ($x_{\rm{H}\,\textsc{ii}}$) bubbles in the IGM. One can certainly expect the DM measurements of the two nearby FRBs to be correlated (Reischke & Hagstotz, 2023) as they traces the underlying structures. To this end, we define the DM structure function, for a given LoS at $\bm{\theta}$ and redshift $z$, as
@@
where $\bm{\delta\theta}$ is a small angular separation for all nearby LoS and $\langle\cdots\rangle_{\bm{\theta},\bm{\widehat{\delta\theta}}}$ denotes double average – first, over different rotations by an amount $\bm{\widehat{\delta\theta}}$ around every $\bm{\theta}$, and next, average over different LoS directions $\bm{\theta}$. This definition utilizes the assumption that the sky is statistically homogeneous and isotropic at any particular $z$, which leaves $\Xi$ a function of the magnitude $\delta\theta$ and $z$ only. The dependence of $\Xi$ on ${\rm{DM}_{EoR}}$, which itself is an integrated quantity, makes it unsuitable to directly conclude anything about the ionized bubble sizes and their growth. Hence, we use the derivatives of $\Xi$ which probe local IGM properties. $\partial\Xi/\partial\delta\theta$ encompasses the information about how fast the structures decorrelates on the sky plane. However it still has integrated effects along the LoS, and we therefore compute the second order derivative $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$. This has both the instantaneous and local information about the structures and their scale information. We will demonstrate how the landscape of $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$ corresponds to the different reionization histories and morphologies in the $(\delta\theta,z)$ plane. Later, we marginalize $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$ over $\delta\theta$ and $z$ one at a time. Marginalization makes it easier to understand the behaviour of $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$ as a function of either $z$ or $\delta\theta$ irrespective of the other variable and requires fewer observed bursts to be determined observationally. We finally compute the average of the derivatives and their marginalized values over various LoS $\bm{\theta}$. This provides us with the information of the mean sizes of the ionized regions in the IGM.
Figure 2: Left: Asymmetrical $\tanh(z)$ ionization histories. Different colors represent three ionization histories corresponding to the simulated toy models. Right: Mean DM estimates and $1\sigma$ fluctuations. Colors represent different reionization histories as shown in the left panel. Solid lines and the shaded regions are the average DM and corresponding $1\sigma$ errors estimated over all grid points available in the simulation volume. Dashed lines correspond to the DM estimated using the mean ionization fraction directly in eq. 3. Three vertical lines mark the redshift of mid reionization.
@@ 3 Toy Model Simulations
We demonstrate the impact of the ionized bubble sizes and the rate of reionization on the estimators mentioned in §2, allowing us to gain intuition and test the general validity of the approach. We use an approximate and simplistic toy model of reionization to simulate the ionization ($x_{\rm{H}\,\textsc{ii}}$) field light-cones (LCs). The toy model assumes all the ionized bubbles have the same radii and that everything inside the bubbles is completely ionized and anything outside is completely neutral. We divide the whole LC boxes into reasonably thin slices along the LoS axis and fill them with a number of spherical bubbles matching the average input ionization fraction for every $z$ slice. We place the bubble centers in the slices uniform randomly and allow overlap between them. A LC thus created will be a binary ionization field where $\Delta(\bm{\theta},z)=1$ and $x_{\rm{H}\,\textsc{ii}}(\bm{\theta},z)$ is either 00 or $1$. This field has basic differences from the LC obtained from the real simulations (see §4.1) where additional fluctuations in the free electron density arise due to perturbations in the underlying matter density field. Furthermore, as opposed to the toy model, the realistic reionization model has inherent temporal growth in the bubble sizes apart from their percolation.
We simulate the toy model LCs (see Figure 1) within a comoving volume $(500\times 500\times 1500)~{}[{h^{-1}~{}\rm Mpc}]^{3}$ that is divided into $(600\times 600\times 1800)$ cubic voxels, accordingly. This particular choice of LC volume is to roughly match with our reionization simulations as described below (§4.1). We use the asymmetric $\tanh(z)$ form of reionization history (e.g., Heinrich et al., 2017; Ghara et al., 2024) for our toy models. This form of the history closely mimics the histories found in simulations. We fix the origin of the toy model LCs at $z=6$ assuming reionization ends by then. The other end of the toy model LC boxes extends up to $z\approx 15$.
@@ 3.1 Dependence on Reionization History
We study the effect of different reionization histories on the ${\rm\overline{{DM}}_{EoR}}$ and the other derived estimators, as defined above. We generate three toy models with ‘Faster’, ‘Fiducial’ and ‘Slower’ reionization histories as shown in Figure 2 with their corresponding DMs. We mimic the ionization histories by varying the number of bubbles in each slice of the LCs with bubble radius being fixed at $10$ grid units ($\approx 12~{}{\rm Mpc}$). We choose the reionization mid point at $z_{\rm mid}=7.5,~{}8.5,~{}9.0$ and the corresponding reionization window to be $\delta z=1.0,~{}2.0,~{}2.5$ (i.e. the reionization to end at the same time for these three toy models) corresponding to the ‘Faster’, ‘Fiducial’ and ‘Slower’ reionization histories, respectively. We fix the asymmetry parameter $\alpha=3$.
@@
As shown in Figure 2, there is a slight offset between ${\rm\overline{{DM}}_{EoR}}$, the mean estimated over all the $600^{2}$ grids on the transverse plane of the box and the average DM estimated using the $\bar{x}_{\rm{H}\,\textsc{ii}}$ into eq. 3 and ignoring the fluctuations along different LoS. As DM is a cumulative estimator, its mean value rises rapidly with $z$ where most of the reionization is happening. At higher $z$ it saturates, as there are no more free electrons to contribute to it. The asymptotic difference between the two mean DM estimators (solid and dashed lines) increases monotonically from faster to slower reionization at any $z$. This happens because the fluctuating structures exists for a larger LoS distance in the case of slower reionization history. We also show the corresponding $1\sigma$ sample variance around ${\rm\overline{{DM}}_{EoR}}$. Fluctuations due to the binary ionization field are not able to cause any significant ($>1\sigma)$ deviations between ${\rm\overline{{DM}}_{EoR}}(z)$ and ${\rm DM}(\bar{x}_{\rm{H}\,\textsc{ii}}(z))$ for all three histories considered here. The deviation should be enhanced for more realistic reionization LC where both $x_{\rm{H}\,\textsc{ii}}$ and $\Delta$ contribute to the LoS fluctuations in ${\rm\overline{{DM}}_{EoR}}$. Also, contribution to $\sigma_{\rm DM}$ accumulated from $z<6$ will increase the spread in ${\rm DM}_{\rm EoR}$.
@@
Figure 3: Top: The ratio $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ as a function of redshift $z$. The three different line styles corresponds to the three reionization histories. Bottom: The redshift derivative of ${\rm\overline{{DM}}_{EoR}}(z)$. The green lines show the fluctuating estimates whereas the orange lines are after a Gaussian smoothing.
Figure 3 shows the ratio $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ that qualitatively follows ${\rm\overline{{DM}}_{EoR}}$ for $z\gtrsim 7$. It first increases rapidly and then saturates towards large $z$ as the ionized regions disappear. However, we find that $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ increases sharply towards the end of reionization ($z<7$). This is because ${\rm\overline{{DM}}_{EoR}}$ has a value around zero at $z\approx 6$ since we ignore the contribution from the low-redshift IGM. Starting at $z\approx 7$, ${\rm\overline{{DM}}_{EoR}}$ increases more rapidly while moving towards higher $z$ than the fluctuations do and finally the ratio saturates. The saturation redshift varies depending on the reionization history. It is at low redshift for the faster reionization and vice-versa. $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ is larger for the faster reionization history and vice-versa, which indicates a direct mapping between the relative (to the mean) fluctuations in the ${\rm DM}_{\rm EoR}$ with the emergence and sizes of the structures in the IGM.
Figure 3 also shows the derivative $d{\rm\overline{{DM}}_{EoR}}/dz$ which directly traces the local electron density. After Gaussian smoothing, the derivatives are roughly the same towards the end of reionization ($z<6.5$) where the IGM has roughly indistinguishable properties. However, the derivatives beyond $z_{\rm mid}$ apparently encodes the information of the reionization history in its slope when plotted against $z$ in a log-log plane. The slower history has a shallower slope and it increases gradually towards fiducial and faster histories. This is simply because there are more ionized bubbles for the slower history causing a larger change in ${\rm\overline{{DM}}_{EoR}}$ at higher $z$.
@@
Figure 4: Density distribution of the double derivative of the structure function $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$. The three different panels correspond to the different reionization histories considered. The different colored filled-contours here represent various percentages of the area encompassed by the $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$ surface as shown in the labels. The horizontal dashed lines mark $z_{\rm mid}$. The dashed lines represent the angular size of bubbles varying with $z$.
[…]
@@
Figure 4 shows contours of the derivatives of the structure function $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$ on the $\delta\theta-z$ plane. We consider only $400$ LoS at every equidistant comoving slice to compute $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$ instead of all available $360,000$ LoS per comoving slice. Our choice of 400400400400111Our mock simulations have access to $400$ FRBs/slice even at larger redshifts. However, in reality, the number of FRBs might drop significantly down at larger $z$ slices (see Figure 15). FRBs per comoving slice is closer to real observations and computationally tractable. On the other hand, we believe that it is a good number for statistical convergence of $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$. We finally bin-average our estimates of all slices within equispaced $z$ bins of width $0.5$ for further use. We consider $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$ as a 2D surface and plot the contours which include top $10\%$, $30\%$, $50\%$ and $67\%$ of the total area under the $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$ surface. For each reionization history we depict $z_{\rm mid}$ and the angular size corresponding to the radius of the individual bubble in our toy model. The peak ($10\%$ contour) gradually shifts to a smaller $z$ while moving from the slower to the faster history, approximately tracking the change in $z_{\rm mid}$, and the contours gets more squeezed along the $z$ axis. This clearly indicates that the peakedness of the $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$ landscape along $z$ is directly connected to the reionization window. There is no considerable change in the extent of the contours along the $\delta\theta$ axis, which is expected since we are using bubbles of fixed radii here.
[…]
@@ 3.2 Dependence on Bubble Sizes
Figure 6: Left: $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ as a function of $z$. The three different linestyles correspond to the different bubble sizes chosen here with radii $R=5,~{}10$ and $20$ (in grid units). Right: Derivative of the ratio shown in left panel w.r.t. $z$. The three different green lines corresponds to the fluctuating estimates from the toy model simulations, whereas the orange lines are after a Gaussian smoothing. The vertical gray line shows the redshift corresponding to $z_{\rm mid}$.
We assess the impact of ionized bubble size on our estimators by fixing the reionization histories to the fiducial case and varying the bubble radius $R$. We consider $R=5,~{}10$ and $20$ grid units which corresponds to $6$, $12$ and $24~{}{\rm Mpc}$, respectively. We have repeated the same analysis as in the §3.1. The three simulations perfectly agree with their common input reionization history, leading to the similar ${\rm\overline{{DM}}_{EoR}}$ values. However, variations in bubble size influence the fluctuations in the ${\rm DM}_{\rm EoR}$ estimates, and consequently $\sigma_{\rm DM}$.
@@
Figure 7 shows the contours of $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$. Similar to Figure 4, we have used $400$ LoS per comoving slices and bin them within a redshift window $\Delta z=0.5$. The contours are nearly unchanged along $z$-axis for the different $R$ values. The peak of the $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$ surface (depicted by $10\%$ contour) shifts towards larger $\delta\theta$ values with increasing $R$, approximately tracking the change in the bubble size. The derivatives decrease for $\delta\theta$ greater than the angular bubble size $\theta_{R}$, as the correlation between the structures decays out. Whereas for scales less than $\theta_{R}$ (see the right panel), the derivatives decrease as the points are tightly correlated and $\Xi$ itself is consistently small.
@@
We can understand this with a simple argument. We distribute the bubbles uniformly in each $z$ slice, following a Poisson distribution while preventing any overlap. Considering a redshift slice, the mean DM would approximately be ${\rm DM}_{1}\times N_{\rm col}$, where ${\rm DM}_{1}$ is the contribution from a single bubble and $N_{\rm col}$ is the average number of bubbles appearing along a LoS. Hence, the corresponding spread in the DM along different LoS would be $\sigma_{\rm DM}\approx{\rm DM}_{1}\times N_{\rm col}^{1/2}$. Since we keep the ionization fraction of slices (and hence mean DM) constant while changing the bubble radius, the total number of the bubbles within the slice varies as $N\propto R^{-3}$. Considering a cubical slice $N_{\rm col}\propto N^{1/3}\propto R^{-1}$. Since ${\rm DM}_{1}\propto R$, we see that $\sigma_{\rm DM}\propto R^{1/2}$. This scaling is consistent with the results plotted in the left panel of Figure 6.
We plot $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ and its $z$ derivative as a function of $z$ in Figure 6. The trend is qualitatively similar to that shown in Figure 3. We find a sharp turnover and rapid rise in $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ by the end of reionization ($z<7$) due to near zero value of ${\rm\overline{{DM}}_{EoR}}$. However for $z>7$, $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ increases almost linearly towards higher $z$ and gradually starts saturating beyond $z\approx 10$. As the saturation point strongly depends on the reionization history, hence the saturation knee is almost at the same redshift for all the three $R$ values. However, the variation in $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ is solely due to variation in $\sigma_{\rm DM}$ with $R$. Overall, $\sigma_{\rm DM}$ increases with increasing $R$ as shown in the Figure. The reason is clear as filling the IGM with the smaller bubbles would make the distribution of the ionized regions roughly uniform and homogeneous, and therefore the fluctuations between the different LoS would be less and vice-versa.
The derivative $d(\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}})/dz$ is a more local quantity as shown in Figure 6. The derivatives drop very sharply approaching $\infty$- ∞ for $z<7$ as there is a rapid increase in $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ with decreasing $z$. However, we see a power-law decrement in the derivative with increasing $z$ for $z\geq 8$. The slope of this power-law drop is roughly the same for the three $R$ values, although the magnitude scales with $R$ in a similar way as for $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$.
[…]

### v2

@@ 2.2 Structure Function
Figure 1: Image slice from the LC box of the fiducial EoR toy model. Both axes are in comoving units, with the axis labeled $x$ being the LoS axis. The bubble radius is $10$ grid units ($\approx 12~{\rm Mpc}$). Black and white regions are, respectively, completely ionized and completely neutral. Reionization progresses from right to left in this box.
Figure 2: Left: asymmetrical $\tanh(z)$ ionization histories. Different colors represent three ionization histories corresponding to the simulated toy models. Right: mean DM estimates and $1\sigma$ fluctuations. Colors represent different reionization histories as shown in the left panel. Solid lines and the shaded regions are the average DM and corresponding $1\sigma$ errors estimated over all grid points available in the simulation volume. Dashed lines correspond to the DM estimated using the mean ionization fraction directly in eq. 3. Three vertical lines mark the redshift of mid-reionization.
We aim to outline how FRBs can be used as a probe for the characteristic size of the ionization ($x_{\rm{H}\,\textsc{ii}}$) bubbles in the IGM. One can certainly expect the DM measurements of the two nearby FRBs to be correlated (Reischke & Hagstotz, 2023) as they trace the underlying structures. To this end, we define the DM structure function, for a given LoS at $\bm{\theta}$ and redshift $z$, as
@@
where $\bm{\delta\theta}$ is a small angular separation for all nearby LoS and $\langle\cdots\rangle_{\bm{\theta},\bm{\widehat{\delta\theta}}}$ denotes double average – first, over different rotations by an amount $\bm{\widehat{\delta\theta}}$ around every $\bm{\theta}$, and next, average over different LoS directions $\bm{\theta}$. This definition utilizes the assumption that the sky is statistically homogeneous and isotropic at any particular $z$, which leaves $\Xi$ a function of the magnitude $\delta\theta$ and $z$ only. The dependence of $\Xi$ on $\rm{DM}_{EoR}$}, which itself is an integrated quantity, makes it unsuitable to directly conclude anything about the ionized bubble sizes and their growth. Hence, we use the derivatives of $\Xi$ which probe local IGM properties. $\partial\Xi/\partial\delta\theta$ encompasses the information about how fast the structures decorrelate on the sky plane. However, it still has integrated effects along the LoS, and we therefore compute the second-order derivative $\partial^{2}\Xi/\partial z~\partial\delta\theta$. This has both the instantaneous and local information about the structures and their scale information. We will demonstrate how the landscape of $\partial^{2}\Xi/\partial z~\partial\delta\theta$ corresponds to the different reionization histories and morphologies in the $(\delta\theta,z)$ plane. Later, we marginalize $\partial^{2}\Xi/\partial z~\partial\delta\theta$ over $\delta\theta$ and $z$ one at a time. Marginalization makes it easier to understand the behavior of $\partial^{2}\Xi/\partial z~\partial\delta\theta$ as a function of either $z$ or $\delta\theta$ irrespective of the other variable, and requires fewer observed bursts to be determined observationally. We finally compute the average of the derivatives and their marginalized values over various LoS $\bm{\theta}$. This provides us with the information on the mean sizes of the ionized regions in the IGM.
@@ 3 Toy Model Simulations
We demonstrate the impact of the ionized bubble sizes and the rate of reionization on the estimators mentioned in §2, allowing us to gain intuition and test the general validity of the approach. We use an approximate and simplistic toy model of reionization to simulate the ionization ($x_{\rm{H}\,\textsc{ii}}$) field LCs. The toy model assumes all the ionized bubbles have the same radius and that everything inside the bubbles is completely ionized and anything outside is completely neutral. We divide the whole LC boxes into reasonably thin slices along the LoS axis and fill them with a number of spherical bubbles matching the average input ionization fraction for every $z$ slice. We place the bubble centers in the slices uniformly at random locations and allow overlap between them. An LC thus created will be a binary ionization field where $\Delta(\bm{\theta},z)=1$ and $x_{\rm{H}\,\textsc{ii}}(\bm{\theta},z)$ is either 0 or $1$. This field has basic differences from the LC obtained from the real simulations (see §4.1), where additional fluctuations in the free electron density arise due to perturbations in the underlying matter density field. Furthermore, as opposed to the toy model, the realistic reionization model has inherent temporal growth [^1] There is no such functional form of the realistic bubble growth with redshift is known in the literature. Using any arbitrary function to model the growth of bubble radii will complicate the results and may not represent a realistic scenario. in the bubble sizes apart from their percolation.
We simulate the toy model LCs (see Figure 1) within a comoving volume $(500\times 500\times 1500)~[{h^{-1}~\rm Mpc}]^{3}$ that is divided into $(600\times 600\times 1800)$ cubic voxels, accordingly. This particular choice of LC volume is to roughly match our reionization simulations as described below (§4.1). We use the asymmetric $\tanh(z)$ form of reionization history (e.g., Heinrich et al., 2017; Ghara, R. et al., 2024) for our toy models. This form of history closely mimics the histories found in simulations. We fix the origin of the toy model LCs at $z=6$, assuming reionization ends by then. The other end of the toy model LC boxes extends up to $z\approx 15$.
@@ 3.1 Dependence on Reionization History
We study the effect of different reionization histories on the $\rm\overline{{DM}}_{EoR}$} and the other derived estimators, as defined above. We generate three toy models with ‘Faster,’ ‘Fiducial,’ and ‘Slower’ reionization histories as shown in Figure 2 with their corresponding DMs. We mimic the ionization histories by varying the number of bubbles in each slice of the LCs, with the bubble radius being fixed at $10$ grid units ($\approx 12~{\rm Mpc}$). We choose the reionization midpoint at $z_{\rm mid}=7.5,~8.5,~9.0$ and the corresponding reionization window to be $\delta z=1.0,~2.0,~2.5$ (i.e. the reionization to end at the same time for these three toy models) corresponding to the ‘Faster,’ ‘Fiducial,’ and ‘Slower’ reionization histories, respectively. We fix the asymmetry parameter $\alpha_{0}=3$ (see eq. 6 of Ghara, R. et al., 2024).
@@
As shown in Figure 2, there is a slight offset between $\rm\overline{{DM}}_{EoR}$}, the mean estimated over all the $600^{2}$ grids on the transverse plane of the box, and the average DM estimated using the $\bar{x}_{\rm{H}\,\textsc{ii}}$ into eq. 3 and ignoring the fluctuations along different LoS. As DM is a cumulative estimator, its mean value rises rapidly with $z$ where most of the reionization is happening. At higher $z$ it saturates, as there are no more free electrons to contribute to it. The asymptotic difference between the two mean DM estimators (solid and dashed lines) increases monotonically from faster to slower reionization at any $z$. This happens because the fluctuating structures exist for a larger LoS distance in the case of slower reionization history. We also show the corresponding $1\sigma$ sample variance around $\rm\overline{{DM}}_{EoR}$}. Fluctuations due to the binary ionization field are not able to cause any significant ($1\sigma)$ deviations between ${\rm\overline{{DM}}_{EoR}}(z)$ and ${\rm DM}(\bar{x}_{\rm{H}\,\textsc{ii}}(z))$ for all three histories considered here. The deviation should be enhanced for more realistic reionization LC where both $x_{\rm{H}\,\textsc{ii}}$ and $\Delta$ contribute to the LoS fluctuations in $\rm\overline{{DM}}_{EoR}$}. Also, contributions to $\sigma_{\rm DM}$ accumulated from $z<6$ will increase the spread in ${\rm DM}_{\rm EoR}$.
@@
Figure 3: Top: the ratio $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ as a function of redshift $z$. The three different line styles correspond to the three reionization histories. Bottom: the redshift derivative of ${\rm\overline{{DM}}_{EoR}}(z)$. The green lines show the fluctuating estimates, whereas the orange lines are after a Gaussian smoothing.
@@
Figure 4: Density distribution of the double derivative of the structure function $\partial^{2}\Xi/\partial z~\partial\delta\theta$. The three different panels correspond to the different reionization histories considered. The different colored filled contours here represent various percentages of the area encompassed by the $\partial^{2}\Xi/\partial z~\partial\delta\theta$ surface, as shown in the labels. The horizontal dashed lines mark $z_{\rm mid}$. The dashed lines represent the angular size of bubbles varying with $z$.
Figure 3 shows the ratio $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ that qualitatively follows $\rm\overline{{DM}}_{EoR}$} for $z\gtrsim 7$. It first increases rapidly and then saturates toward large $z$ as the ionized regions disappear. However, we find that $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ increases sharply toward the end of reionization ($z<7$). This is because $\rm\overline{{DM}}_{EoR}$} has a value around zero at $z\approx 6$ since we do not consider the contribution from the low-redshift IGM. Starting at $z\approx 7$, $\rm\overline{{DM}}_{EoR}$} increases more rapidly while moving toward higher $z$ than the fluctuations do, and finally the ratio saturates. The saturation redshift varies depending on the reionization history. It is at low redshift for the faster reionization and vice versa. $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ is larger for the faster reionization history and vice versa, which indicates a direct mapping between the relative (to the mean) fluctuations in the ${\rm DM}_{\rm EoR}$ with the emergence and sizes of the structures in the IGM.
Figure 3 also shows the derivative $d{\rm\overline{{DM}}_{EoR}}/dz$ which directly traces the local electron density. After Gaussian smoothing, the derivatives are roughly the same toward the end of reionization ($z<6.5$) where the IGM has roughly indistinguishable properties. However, the derivatives beyond $z_{\rm mid}$ apparently encode the information of the reionization history in its slope when plotted against $z$ in a log-log plane. The slower history has a shallower slope, and it increases gradually toward fiducial and faster histories. This is simply because there are more ionized bubbles for the slower history, causing a larger change in $\rm\overline{{DM}}_{EoR}$} at higher $z$.
Figure 5: The distribution shown in Figure 4 marginalized along $\delta\theta$. The three solid lines represent $\int(\partial^{2}\Xi/\partial z~\partial\delta\theta)~d\delta\theta$ for the three reionization histories. The dashed vertical lines correspond to the respective $z_{\rm mid}$ values.
Figure 6: Left: $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ as a function of $z$. The three different line styles correspond to the different bubble sizes chosen here with radii R=5,10R=5,~10 and 2020 (in grid units). Right: derivative of the ratio shown in the left panel with respect to $z$. The three different green lines correspond to the fluctuating estimates from the toy model simulations, whereas the orange lines are after a Gaussian smoothing. The vertical gray line shows the redshift corresponding to $z_{\rm mid}$.
@@
Figure 4 shows contours of the derivatives of the structure function $\partial^{2}\Xi/\partial z~\partial\delta\theta$ on the $\delta\theta-z$ plane. We randomly consider only $400$ LoS at every equidistant comoving slice to compute $\partial^{2}\Xi/\partial z~\partial\delta\theta$ instead of all available $360,000$ LoS per comoving slice. Our choice of 400400222Our mock simulations have access to $400$ FRBs/slice even at larger redshifts. However, in reality, the number of FRBs might drop significantly at larger $z$ slices (see Figure 15). FRBs per comoving slice is closer to real observations and computationally tractable. On the other hand, we believe that it is a good number for statistical convergence of $\partial^{2}\Xi/\partial z~\partial\delta\theta$ as it does not change much when choosing different sets of $400$ LoS. We finally bin-average our estimates of all slices within equispaced $z$ bins of width $0.5$ for further use. Our choice of bin-width is a trade-off between capturing the evolution of the estimators and getting sufficiently good statistics per bin. We consider $\partial^{2}\Xi/\partial z~\partial\delta\theta$ as a 2D surface and plot the contours, which include the top $10\%$, $30\%$, $50\%$ and $67\%$ of the total area under the $\partial^{2}\Xi/\partial z~\partial\delta\theta$ surface. For each reionization history, we depict $z_{\rm mid}$ and the angular size corresponding to the radius of the individual bubble in our toy model. The peak ($10\%$ contour) gradually shifts to a smaller $z$ while moving from the slower to the faster history, approximately tracking the change in $z_{\rm mid}$, and the contours get more squeezed along the $z$- axis. This clearly indicates that the peakedness of the $\partial^{2}\Xi/\partial z~\partial\delta\theta$ landscape along $z$ is directly connected to the reionization window. There is no considerable change in the extent of the contours along the $\delta\theta$ axis, which is expected since we are using bubbles of fixed radii here.
We next marginalize $\partial^{2}\Xi/\partial z~\partial\delta\theta$ over $\delta\theta$. The marginalized result, $\int(\partial^{2}\Xi/\partial z~\partial\delta\theta)~d{\delta\theta}$, is shown in Figure 5 as a function of $z$. $\int(\partial^{2}\Xi/\partial z~\partial\delta\theta)~d{\delta\theta}$ peaks roughly around $z_{\rm mid}$. The decrease toward the end of EoR is sharper and roughly independent of reionization history. However, the drop toward higher $z$ is related to the reionization history. The drop is slower for the slower reionization history and vice versa. That could be simply related to the rate of emergence of the ionized bubbles in the IGM, and once the reionization is roughly around its midway, the percolation of bubbles makes it insensitive to the history.
@@ 3.2 Dependence on Bubble Sizes
We assess the impact of ionized bubble size on our estimators by fixing the reionization histories to the fiducial case and varying the bubble radius $R$. We consider R=5,10R=5,~10 and $20$ grid units, which correspond to $6$, $12$, and $24~{\rm Mpc}$, respectively. We have repeated the same analysis as in §3.1. The three simulations perfectly agree with their common input reionization history, leading to similar $\rm\overline{{DM}}_{EoR}$} values. However, variations in bubble size influence the fluctuations in the ${\rm DM}_{\rm EoR}$ estimates, and consequently, $\sigma_{\rm DM}$.
@@
We plot $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ and its $z$ derivative as a function of $z$ in Figure 6. The trend is qualitatively similar to that shown in Figure 3. We find a sharp turnover and rapid rise in $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ by the end of reionization ($z<7$) due to a near-zero value of $\rm\overline{{DM}}_{EoR}$}. However, for $z>7$, $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ increases almost linearly toward higher $z$ and gradually starts saturating beyond $z\approx 10$. As the saturation point strongly depends on the reionization history, the saturation knee is almost at the same redshift for all three $R$ values. However, the variation in $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ is solely due to variation in $\sigma_{\rm DM}$ with $R$. Overall, $\sigma_{\rm DM}$ increases with increasing $R$ as shown in the Figure. The reason is clear, as filling the IGM with the smaller bubbles would make the distribution of the ionized regions roughly uniform and homogeneous, and therefore the fluctuations between the different LoS would be less, and vice versa.
@@
We can understand this with a simple argument. We distribute the bubbles uniformly in each $z$ slice, following a Poisson distribution while preventing any overlap. Considering a redshift slice, the mean DM would approximately be ${\rm DM}_{1}\times N_{\rm col}$, where ${\rm DM}_{1}$ is the contribution from a single bubble and $N_{\rm col}$ is the average number of bubbles appearing along a LoS. Hence, the corresponding spread in the DM along different LoS would be $\sigma_{\rm DM}\approx{\rm DM}_{1}\times N_{\rm col}^{1/2}$. Since we keep the ionization fraction of slices (and hence mean DM) constant while changing the bubble radius, the total number of bubbles within the slice varies as $N\propto R^{-3}$. Considering a cubical slice $N_{\rm col}\propto N^{1/3}\propto R^{-1}$. Since ${\rm DM}_{1}\propto R$, we see that $\sigma_{\rm DM}\propto R^{1/2}$. This scaling is consistent with the results plotted in the left panel of Figure 6.
The derivative $d(\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}})/dz$ is a more local quantity, as shown in Figure 6. The derivatives drop very sharply approaching -$\infty$ for $z<7$ as there is a rapid increase in $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ with decreasing $z$. However, we see a power-law decrement in the derivative with increasing $z$ for $z\geq 8$. The slope of this power-law drop is roughly the same for the three $R$ values, although the magnitude scales with $R$ in a similar way as for $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$.
Figure 7 shows the contours of $\partial^{2}\Xi/\partial z~\partial\delta\theta$. Similar to Figure 4, we have used $400$ LoS per comoving slices and bin them within a redshift window $\Delta z=0.5$. The contours are nearly unchanged along the $z$-axis for the different $R$ values. The peak of the $\partial^{2}\Xi/\partial z~\partial\delta\theta$ surface (depicted by a $10\%$ contour) shifts toward larger $\delta\theta$ values with increasing $R$, approximately tracking the change in the bubble size. The derivatives decrease for $\delta\theta$ greater than the angular bubble size $\theta_{R}$, as the correlation between the structures decays out. Whereas for scales less than $\theta_{R}$ (see the right panel), the derivatives decrease as the points are tightly correlated and $\Xi$ itself is consistently small.
We again marginalize $\partial^{2}\Xi/\partial z~\partial\delta\theta$ here, but this time along the $z$-axis, to obtain $\int(\partial^{2}\Xi/\partial z~\partial\delta\theta)~dz$ as shown in Figure 8 for the three $R$ values. We observe a similar qualitative trend as seen in the peaks of the contour plots. The peak of $\int(\partial^{2}\Xi/\partial z~\partial\delta\theta)~dz$ shifts to a larger $\delta\theta$ for large bubble sizes. We also note that it decreases for $\delta\theta$ larger than $\theta_{R}$. $\langle\theta_{R}\rangle$ is computed by marginalizing the $\theta_{R}(z)$ corresponding to the respective characteristic bubble size $R$ in the $z$ range of interest. Here, the sharp fall in the curves with $R$ justifies that the angular correlations drop rapidly at scales larger than the bubble radii.